Lithium has two naturally occurring isotopes: lithium-6 and lithium-7. If the average atomic mass of lithium is 6.941 atomic mass units, which isotope is the most abundant? How do you know?
1
Expert's answer
2020-12-03T13:50:30-0500

Let "w(isotope)" stand for the ratio of the overall amount of that isotope to the overall amount of all isotopes of the specified element.


Then for "Li" the following holds:

"w(^6Li) + w(^7Li) = 1"

and


"6*w(^6Li) + 7*w(^7Li) = 6.941"


From the first equation:

"w(^6Li) = 1 - w(^7Li)"


From this and the second equation:

"6*(1 - w(^7Li)) + 7*w(^7Li) = 6.941"

"6 - 6* w(^7Li) + 7*w(^7Li) = 6.941"

"w(^7Li) = 0.941"

"w(^6Li) = 0.059"


"w(^6Li) < w(^7Li)" , therefore "^7Li" is the most abundant isotope.


Answer: "^7Li" is the most abundant isotope